Use the discriminate to determine the number of and type of solutions.
-3m2 - 7m = 6

A. 2 rational solutions

B. 2 irrational solutions

C. 1 irrational solution

D. 0 solutions

Answer by Alan3354(69443)

Use the discriminate to determine the number of and type of solutions.
-3m2 - 7m = 6
3m^2 + 7m + 6 = 0
Use ^ for exponents.
-----------
Discriminant D = b^2 - 4ac
------
There are always 2 solutions.
If D = 0, the 2 solutions are equal and rational.
-----
If D is a perfect square > 0, there are 2 rational solutions.
If D < 0 there are 2 complex solutions (a conjugate pair).
-----
PS Discriminant, not discriminate.